# A Mixed-Integer Conic Program for the Moving-Target Traveling Salesman Problem based on a Graph of Convex Sets

### v1

@@ IV-A Biconvex Formulation
constraints (2), (3), (4), (5), (6)
$p_{i,x}=\underline{p}_{i,x}+t_{i}v_{i,x}-\underline{t}_{i}v_{i,x},\;\;\forall\;i\in V$ (13)
$p_{i,y}=\underline{p}_{i,y}+t_{i}v_{i,y}-\underline{t}_{i}v_{i,y},\;\;\forall\;i\in V$ (14)
$l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=(z^{\prime}_{e,x}-z_{e,x}),\;\;\forall e=(i,j)\in E$ (15)
$l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=(z^{\prime}_{e,y}-z_{e,y}),\;\;\forall e=(i,j)\in E$ (16)
$l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\leq v_{max}(z^{\prime}_{e,t}-z_{e,t}),\;\;\forall e=(i,j)\in E$ (17)
$\begin{split}&(l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e}))^{2}+(l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e}))^{2}\leq(l(\tilde{z}_{e},\tilde{z}^{\prime}_{e}))^{2},\\ &\forall\;e=(i,j)\in E\end{split}$ (18)
$z_{e}=(y_{e}p_{i},y_{e}t_{i}),\;z^{\prime}_{e}=(y_{e}p_{j},y_{e}t_{j}),\;\;\forall e=(i,j)\in E$ (19)
This formulation shares constraints (2) to (6) from the MICP. Constraints (13) and (14) describes variables $p_{i,x}$ and $p_{i,y}$. The nonconvexity of this program comes only from the bilinear constraints, (19). These constraints will be utilized to achieve the role satisfied by the big-$M$ constraints in the MICP. First, notice how for each edge $e=(i,j)\in E$, when $y_{e}=0$, (19) becomes $z_{e}=z^{\prime}_{e}=(0,0,0)$. Consequently, (15), (16) becomes $l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=0$. However, when $y_{e}=1$, (19) becomes $z_{e}=(p_{i},t_{i})$, $z^{\prime}_{e}=(p_{j},t_{j})$ and consequently, (15), (16) becomes $l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=p_{j,x}-p_{i,x}$, and $l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=p_{j,y}-p_{i,y}$.
Now, consider (17). For each edge $e=(i,j)\in E$, when $y_{e}=1$, these constraints become $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\leq v_{max}(t_{j}-t_{i})$. However, when $y_{e}=0$, we get $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\leq 0$, allowing $t_{i}$, and $t_{j}$ to be free. Finally, consider (18). We see how these constraints with the help of (15), (16) becomes $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\geq\|p_{j}-p_{i}\|$ when $y_{e}=1$, but allows $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})$ to take any value when $y_{e}=0$. Recall how the big-$M$ constraints, (9), (10), and the constraints, (11) achieves the same role as (17) and (18) for $\tilde{l}(p_{i},p_{j})$. Hence, in the biconvex formulation, constraints (17) and (18) replaces constraints (9), (10) and (11) from the MICP, while establishing the relationship,
$l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=\tilde{l}(p_{i},p_{j})$ (20)
From (20), we see how both the MICP and the biconvex formulation have the same optimal value. Moreover, we can recover an optimal tour for the MT-TSP from the solution of the biconvex program by picking all the edges with $y_{e}=1$, and recovering the $(p_{i},t_{i})$ for each node $i\in V$.
@@
Now, we discuss the key idea behind the GCS-based MICP we introduce in this paper. Notice how the set of all $(p_{i,x},p_{i,y},t_{i})$ that satisfy (6), (13), and (14) for each node $i\in V$ represents the trajectory-segment of that node within its time-window. These trajectory-segments are line segments when expressed within the $(x,y,t)$ coordinate system as shown in the example illustration in Fig. 2. Hence, the trajectory-segment corresponding to each node $i$ can be considered a convex set $X_{i}$ corresponding to that node, and the agent is required to visit a point within $X_{i}$. This allows us to solve the MT-TSP by leveraging the ideas presented in [28], where the problem was to find the shortest paths in graphs of convex sets. To summarize, (6), (13), and (14), together can be represented using the following set of constraints.
$(p_{i},t_{i})\in X_{i},\;\;\forall\;i\in V$ (21)
Note that the nonconvexities which arise from the bilinear constraints makes solving the biconvex formulation very challenging. However in the next section, we will introduce our MICP-GCS formulation for the MT-TSP, that can be easily handled by standard solvers.
@@ IV-B GCS-Based Mixed Integer Conic Program (MICP-GCS)
constraints (2), (3), (4), (15), (16), (17), (18)
$\sum_{e\in E^{in}_{i}}(z^{\prime}_{e,t},y_{e})=\sum_{e\in E^{out}_{i}}(z_{e,t},y_{e}),\;\;\forall\;i\in V_{tar}$ (23)
$(z_{e},y_{e})\in\tilde{X}_{i},\;(z^{\prime}_{e},y_{e})\in\tilde{X}_{j},\;\;\forall\;e=(i,j)\in E$ (24)
The MICP-GCS formulation differs from the the biconvex formulation in the following ways: The constraints, (23) are obtained by combining (5) with the additional constraints, $\sum_{e\in E^{in}_{i}}z^{\prime}_{e,t}=\sum_{e\in E^{out}_{i}}z_{e,t}$ for each node $i\in V_{tar}$. These additional constraints ensures that the time at which the agent visits each target $i$ will be equal to the time at which the agent departs from target $i$. Also, (21) which encapsulates (6), (13), (14), as well as the bilinear constraints, (19) are replaced by the constraints, (24). These constraints requires that for each edge $e=(i,j)\in E$, $(z_{e},y_{e})$ and $(z^{\prime}_{e},y_{e})$ lies within the perspective of the convex sets $X_{i}$ and $X_{j}$ respectively. This is a compact way of representing the set of constraints for all the edges $e=(i,j)\in E$ as below:
$y_{e}\underline{t}_{i}\leq z_{e,t}\leq y_{e}\overline{t}_{i}$ (25)
$y_{e}\underline{t}_{j}\leq z^{\prime}_{e,t}\leq y_{e}\overline{t}_{j}$ (26)
$z_{e,x}-v_{i,x}z_{e,t}-y_{e}(\underline{p}_{i,x}-\underline{t}_{i}v_{i,x})=0$ (27)
$z_{e,y}-v_{i,y}z_{e,t}-y_{e}(\underline{p}_{i,y}-\underline{t}_{i}v_{i,y})=0$ (28)
$z^{\prime}_{e,x}-v_{j,x}z^{\prime}_{e,t}-y_{e}(\underline{p}_{j,x}-\underline{t}_{j}v_{j,x})=0$ (29)
$z^{\prime}_{e,y}-v_{j,y}z^{\prime}_{e,t}-y_{e}(\underline{p}_{j,y}-\underline{t}_{j}v_{j,y})=0$ (30)
From an optimal solution to the MICP-GCS, we can recover an optimal agent tour for the biconvex program as follows: For each node $i\in V$, find the optimal $(p_{i},t_{i})$, using the following equations:
$(p_{i},t_{i})=\sum_{e\in E^{in}_{i}}z^{\prime}_{e},\;\;\forall\;i\in V\setminus\{s\}$ (31)
$(p_{s},t_{s})=\sum_{e\in E^{out}_{s}}z_{e}$ (32)
@@ IV-C Proof of Validity
Let $E_{tour}$ be the set of all edges with $y_{e}=1$, obtained from a solution to either of the two formulations. The constraints in both the formulations requires that the edges in $E_{tour}$ forms a path that starts at $s$, visits all the target nodes once, and ends at $s^{\prime}$, thereby forming an agent tour. For each edge $e\notin E_{tour}$, $z_{e}=z^{\prime}_{e}=(0,0,0)$ for both the formulations. This is achieved by (19) in the biconvex formulation, and by (24) in the MICP-GCS formulation. Consequently, the cost addends corresponding to these edges becomes $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=0$ for both formulations. Now, consider each edge $e=(i,j)\in E_{tour}$. In the biconvex formulation, (19) becomes $z_{e}=(p_{i},t_{i})$, $z^{\prime}_{e}=(p_{j},t_{j})$. Additionally, (21) requires $(p_{i},t_{i})\in X_{i}$, $(p_{j},t_{j})\in X_{j}$. These two requirements are the same as saying $z_{e}\in X_{i}$, $z^{\prime}_{e}\in X_{j}$, and for any two adjacent edges $e=(i,j)$ and $f=(j,k)$ in the agent tour, $z^{\prime}_{e}=z_{f}$. In the MICP-GCS formulation, (24) becomes $z_{e}\in X_{i}$, $z^{\prime}_{e}\in X_{j}$, and the additional flow requirement in (23) ensures that for any two adjacent edges $e=(i,j)$ and $f=(j,k)$ in the agent tour, $z^{\prime}_{e}=z_{f}$. Therefore, the cost addends corresponding to edges $e\in E_{tour}$ becomes $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})$ for both the formulations. Now, suppose that we have a solution to the MICP-GCS formulation. The $(p_{i},t_{i})$ corresponding to each $i\in V$ can then be obtained as shown in (31) and (32) since it ensures $(p_{i},t_{i})=z_{e}$ and $(p_{j},t_{j})=z^{\prime}_{e}$ for each $e=(i,j)\in E$ with $y_{e}=1$. ∎
@@ V-A Test Settings and Instance Generation
All the tests were run on a laptop with an Intel Core I7-7700HQ 2.80GHz CPU, and 16GB RAM. The implementation was in Python 3.11.6, and both the MICP as well as MICP-GCS formulations were solved using Gurobi 10.0.3 optimizer. All the Gurobi parameters were set to their default values, except for TimeLimit [^1] Limits the total time expended (in seconds)., which was set to 1800.
@@
A total of 80 instances were generated, 20 each for 5, 10, 15, and 20 targets. The instances were defined by the number of targets $n$, a square area of fixed size $S=100$ units with corresponding diagonal length $R=\sqrt{2}S$, a fixed time-horizon $T=150$, the depot location fixed at the center $(0,0)$ of the square, and finally, a set of randomly generated linear trajectories corresponding to the $n$ targets such that each target has a constant speed within $[0.5,1]$ and is confined within the square area.
@@
For each instance, we ran experiments where we varied two additional test parameters: the max agent speed $v_{max}$, and the time-windows corresponding to the targets. $v_{max}$ was varied to be 4, 6, and 8 $unit/sec$, and the time-windows were varied to be of durations 25, 50, and 75 secs. The time-windows were selected such that a feasible solution could be found for all the generated instances, with all agent speed choices. To do this, we first randomly chose a sequence in which the agent visits the targets, and then found the quickest agent tour corresponding to that sequence by fixing the agent speed at its lowest choice (which is 4 $unit/sec$). If the time for the tour was more than $T$, we tried another random sequence. Otherwise, we took the times where the agent visited each target, and defined time-windows that contained these times. Note that when varying time-windows, we ensured that the time-window of duration 25 lies within the time-window of duration 50, which then lies within the time-window of duration 75, for each target.
V-B Evaluating the Formulations
@@
Given $v_{max}$, a time-window duration, and the formulation of choice, the solver is first run on all the 20 instances corresponding to a given number of targets. The optimality gap value from the solver for an instance is defined as $\frac{|z_{P}-z_{D}|}{|z_{P}|}\times 100$, where $z_{P}$ is the primal (feasible) objective, and $z_{D}$ is the dual (lower-bound) objective. Let % Gap denote the average of the smallest gap values output by the solver for all these instances, and let runtime be the average of the run-times output by the solver for all these instances.
V-C Varying the Time-Window Duration

### v3

@@ IV-A Biconvex Binary Program for the MT-TSP
subject to constraints (2), (3), (4), (5), (6),
$p_{i,x}=\underline{p}_{i,x}+t_{i}v_{i,x}-\underline{t}_{i}v_{i,x},\;\;\forall\;i\in V,$ (13)
$p_{i,y}=\underline{p}_{i,y}+t_{i}v_{i,y}-\underline{t}_{i}v_{i,y},\;\;\forall\;i\in V,$ (14)
$l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=(z^{\prime}_{e,x}-z_{e,x}),\;\;\forall e=(i,j)\in E,$ (15)
$l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=(z^{\prime}_{e,y}-z_{e,y}),\;\;\forall e=(i,j)\in E,$ (16)
$l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\leq v_{max}(z^{\prime}_{e,t}-z_{e,t}),\;\;\forall e=(i,j)\in E,$ (17)
$\begin{split}&(l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e}))^{2}+(l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e}))^{2}\leq(l(\tilde{z}_{e},\tilde{z}^{\prime}_{e}))^{2},\\ &\forall\;e=(i,j)\in E,\end{split}$ (18)
$z_{e}=(y_{e}p_{i},y_{e}t_{i}),\;z^{\prime}_{e}=(y_{e}p_{j},y_{e}t_{j}),\;\;\forall e=(i,j)\in E.$ (19)
This formulation shares constraints (2) to (6) from the MICP. Constraints (13) and (14) describe variables $p_{i,x}$ and $p_{i,y}$. Apart from the binary requirement for all the $y_{e}$ variables, the non-convexities of this program come only from the bilinear constraints, (19). These constraints will be utilized to achieve the role satisfied by the big-$M$ constraints in the MICP. First, notice how for each edge $e=(i,j)\in E$, when $y_{e}=0$, (19) becomes $z_{e}=z^{\prime}_{e}=(0,0,0)$. Consequently, (15), (16) become $l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=0$. However, when $y_{e}=1$, (19) becomes $z_{e}=(p_{i},t_{i})$, $z^{\prime}_{e}=(p_{j},t_{j})$ and consequently, (15), (16) become $l_{x}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=p_{j,x}-p_{i,x}$, and $l_{y}(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=p_{j,y}-p_{i,y}$.
Now, consider (17). For each edge $e=(i,j)\in E$, when $y_{e}=1$, these constraints become $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\leq v_{max}(t_{j}-t_{i})$. However, when $y_{e}=0$, we get $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\leq 0$, allowing $t_{i}$, and $t_{j}$ to be free. Finally, consider (18). We see how these constraints with the help of (15), (16) become $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})\geq\left\lVert p_{j}-p_{i}\right\rVert_{2}$ when $y_{e}=1$, but allow $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})$ to take any value when $y_{e}=0$. Recall how the big-$M$ constraints, (9), (10), and the constraints, (11) achieve the same role as (17) and (18) for $\tilde{l}(p_{i},p_{j})$. Hence, in the biconvex formulation, constraints (17) and (18) replace constraints (9), (10) and (11) from the MICP, while establishing the relationship,
$l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=\tilde{l}(p_{i},p_{j}).$ (20)
From (20), we see how both the MICP and the biconvex formulation have the same optimal value. Moreover, we can recover an optimal tour for the MT-TSP from the solution of the biconvex formulation by recovering the $(p_{i},t_{i})$ for each node $i\in V$.
@@
Now, we discuss the key idea behind the GCS-based MICP we introduce in this paper. Notice how the set of all $(p_{i,x},p_{i,y},t_{i})$ that satisfy (6), (13), and (14) for each node $i\in V$ represents the trajectory-segment of that node within its time-window. These trajectory-segments are line segments when expressed within the $(x,y,t)$ coordinate system as shown in the example illustration in Fig. 2. Hence, the trajectory-segment corresponding to each node $i$ can be considered a convex set $X_{i}$ corresponding to that node, and the agent is required to visit a point within $X_{i}$. This allows us to solve the MT-TSP by leveraging the ideas presented in [28], where the problem was to find the shortest paths in graphs of convex sets. To summarize, (6), (13), and (14), together can be represented using the following set of constraints:
$(p_{i},t_{i})\in X_{i},\;\;\forall\;i\in V.$ (21)
Note that the non-convexities that arise from the bilinear constraints make the biconvex formulation very challenging to solve. However, in the next section, we will introduce our MICP-GCS formulation for the MT-TSP, which can be easily handled by standard solvers.
@@ IV-B GCS-Based Mixed Integer Conic Program (MICP-GCS)
subject to constraints (2), (3), (4), (15), (16), (17), (18),
$\sum_{e\in E^{in}_{i}}(z^{\prime}_{e,t},y_{e})=\sum_{e\in E^{out}_{i}}(z_{e,t},y_{e}),\;\;\forall\;i\in V_{tar},$ (23)
$(z_{e},y_{e})\in\tilde{X}_{i},\;(z^{\prime}_{e},y_{e})\in\tilde{X}_{j},\;\;\forall\;e=(i,j)\in E.$ (24)
The MICP-GCS formulation differs from the biconvex formulation in the following ways: The constraints, (23) are obtained by combining (5) with the additional constraints, $\sum_{e\in E^{in}_{i}}z^{\prime}_{e,t}=\sum_{e\in E^{out}_{i}}z_{e,t}$ for each node $i\in V_{tar}$. These additional constraints ensure that the time at which the agent visits each target $i$ will be equal to the time at which the agent departs from target $i$. Additionally, (21) which encapsulates (6), (13), (14), as well as the bilinear constraints, (19) are replaced by the constraints, (24). These constraints require that for each edge $e=(i,j)\in E$, $(z_{e},y_{e})$ and $(z^{\prime}_{e},y_{e})$ lie within the perspective [^1] The perspective of a compact, convex set $X\subset\mathbb{R}^{n}$ is defined as $\tilde{X}:=\{(x,\lambda):\lambda\geq 0,x\in\lambda X\}$. of the convex sets $X_{i}$ and $X_{j}$ respectively. This is a compact way of representing the set of constraints for all the edges $e=(i,j)\in E$ as shown below:
$y_{e}\underline{t}_{i}\leq z_{e,t}\leq y_{e}\overline{t}_{i},$ (25)
$y_{e}\underline{t}_{j}\leq z^{\prime}_{e,t}\leq y_{e}\overline{t}_{j},$ (26)
$z_{e,x}-v_{i,x}z_{e,t}-y_{e}(\underline{p}_{i,x}-\underline{t}_{i}v_{i,x})=0,$ (27)
$z_{e,y}-v_{i,y}z_{e,t}-y_{e}(\underline{p}_{i,y}-\underline{t}_{i}v_{i,y})=0,$ (28)
$z^{\prime}_{e,x}-v_{j,x}z^{\prime}_{e,t}-y_{e}(\underline{p}_{j,x}-\underline{t}_{j}v_{j,x})=0,$ (29)
$z^{\prime}_{e,y}-v_{j,y}z^{\prime}_{e,t}-y_{e}(\underline{p}_{j,y}-\underline{t}_{j}v_{j,y})=0.$ (30)
From an optimal solution to the MICP-GCS, we can recover an optimal agent tour for the biconvex formulation as follows: For each node $i\in V$, find the optimal $(p_{i},t_{i})$, using the following equations:
$(p_{i},t_{i})=\sum_{e\in E^{in}_{i}}z^{\prime}_{e},\;\;\forall\;i\in V\setminus\{s\},$ (31)
$(p_{s},t_{s})=\sum_{e\in E^{out}_{s}}z_{e}.$ (32)
@@ IV-C Proof of Validity
Let $E_{tour}$ be the set of all edges with $y_{e}=1$, obtained from a solution to either of the two formulations. The constraints in both the formulations require that the edges in $E_{tour}$ form a path that starts at $s$, visits all the target nodes once, and ends at $s^{\prime}$, thereby forming an agent tour. For each edge $e\notin E_{tour}$, $z_{e}=z^{\prime}_{e}=(0,0,0)$ for both the formulations. This is achieved by (19) in the biconvex formulation, and by (24) in the MICP-GCS formulation. Consequently, the cost addends corresponding to these edges become $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})=0$ for both formulations. Now, consider each edge $e=(i,j)\in E_{tour}$. In the biconvex formulation, (19) becomes $z_{e}=(p_{i},t_{i})$, $z^{\prime}_{e}=(p_{j},t_{j})$. Additionally, (21) requires $(p_{i},t_{i})\in X_{i}$, $(p_{j},t_{j})\in X_{j}$. These two requirements are the same as saying $z_{e}\in X_{i}$, $z^{\prime}_{e}\in X_{j}$, and for any two adjacent edges $e=(i,j)$ and $f=(j,k)$ in the agent tour, $z^{\prime}_{e}=z_{f}$. In the MICP-GCS formulation, (24) becomes $z_{e}\in X_{i}$, $z^{\prime}_{e}\in X_{j}$, and the additional flow requirement in (23) ensures that for any two adjacent edges $e=(i,j)$ and $f=(j,k)$ in the agent tour, $z^{\prime}_{e}=z_{f}$. Therefore, the cost addends corresponding to edges $e\in E_{tour}$ become $l(\tilde{z}_{e},\tilde{z}^{\prime}_{e})$ for both the formulations. Now, suppose that we have a solution to the MICP-GCS formulation. The $(p_{i},t_{i})$ corresponding to each $i\in V$ can then be obtained as shown in (31) and (32) since they ensure $(p_{i},t_{i})=z_{e}$ and $(p_{j},t_{j})=z^{\prime}_{e}$ for each $e=(i,j)\in E$ with $y_{e}=1$.
@@ V-A Test Settings and Instance Generation
All the tests were run on a laptop with an Intel Core I7-7700HQ 2.80GHz CPU, and 16GB RAM. The implementation was in Python 3.11.6, and both the MICP as well as MICP-GCS formulations were solved using Gurobi 10.0.3 optimizer [29]. All the Gurobi parameters were set to their default values, except for TimeLimit [^2] Limits the total time expended (in seconds)., which was set to 1800.
@@
A total of 80 instances were generated, 20 each for 5, 10, 15, and 20 targets. The instances were defined by the number of targets $n$, a square area of fixed size $S=100$ units with corresponding diagonal length $R=\sqrt{2}S$, a fixed time-horizon $T=150$ secs, the depot location fixed at the center $(0,0)$ of the square, and finally, a set of randomly generated linear trajectories corresponding to the $n$ targets such that each target has a constant speed within $[0.5,1]$ $unit/sec$ and is confined within the square area.
@@
For each instance, we ran experiments where we varied two additional test parameters: the maximum agent speed $v_{max}$, and the time-windows corresponding to the targets. $v_{max}$ was varied to be 4, 6, and 8 $unit/sec$, and the time-windows were varied to be of durations 25, 50, and 75 secs. The time-windows were selected such that a feasible solution could be found for all the generated instances, with all $v_{max}$ choices. To do this, we first randomly chose a sequence in which the agent visits the targets, and then found the quickest agent tour corresponding to that sequence by fixing $v_{max}$ at its lowest choice (which is 4 $unit/sec$). If the time for the tour was more than $T$, we tried another random sequence. Otherwise, we took the times when the agent visited each target, and defined time-windows that contained these times. Note that when varying time-windows, we ensured that the time-window of duration 25 lies within the time-window of duration 50, which then lies within the time-window of duration 75, for each target.
@@
To evaluate the MICP and MICP-GCS formulations, we use % Gap, and runtime, which we will now explain. Given $v_{max}$, a time-window duration, and the formulation of choice, the solver is first run on all the 20 instances corresponding to a given number of targets. The optimality gap value from the solver for an instance is defined as $\frac{|z_{P}-z_{D}|}{|z_{P}|}\times 100$, where $z_{P}$ is the primal (feasible) objective, and $z_{D}$ is the dual (lower-bound) objective. % Gap denotes the average of the smallest gap values output by the solver for all these instances, and runtime denotes the average of the run-times output by the solver for all these instances.
V-B Varying the Time-Window Duration